You are given 10 mL of a 6× stock solution of 1 M NaCl (MW 58.443), and told to dilute it 6-fold. What is the molarity of this diluted NaCl solution?
Answer #1

use dilution formula
M1*V1 = M2*V2
1---> is for stock solution
2---> is for diluted solution

Given:
M1 = 1 M
V1 = 10 mL
V2 = 60 mL [Diluting it 6 fold means final volume is 6 times initial volume]

use:
M1*V1 = M2*V2
M2 = (M1*V1)/V2
M2 = (1*10)/60
M2 = 0.1667 M
Answer: 0.167 M